OpenStudy (anonymous):

find three consecutive even numbers so that the sum of the first and third numbers is 22 less than then three times the second number

OpenStudy (anonymous):

First number = x Second number = x+2 Third number = x+4 \[(x)+(x+4) = 3(x+2)-22\] \[2x + 4 = 3x+6 - 22\]\[-x+4 = -16\]\[x = 20\]

OpenStudy (anonymous):

So 20, 22, 24

OpenStudy (anonymous):

thanks!